OpenStudy (anonymous):

atch the lengths of the hypotenuse and one leg of a right triangle to the length of the other leg. Hypotenuse: 41; Leg: 40 A) Leg: 5 Hypotenuse: 6; Leg: sqr rt of27 B) Leg: 8 Hypotenuse: sqr rt of 29 ; Leg: 2 C) Leg: 3 Hypotenuse: 10; Leg: 6 D) Leg: 9

OpenStudy (anonymous):

just use the pythagorean theorem for example since a^2+b^2=c^2 40^2+b^2=41^2 1600+b^2=1681 b^2=81 b=9 so on

OpenStudy (anonymous):

D) C) A) B)

OpenStudy (anonymous):

U ROCK!!! lmaoo :D